# Adaptive quantum dynamics with the time-dependent variational Monte Carlo method

### v3

@@ II.3.3 Collective parameter update
For situations that demand a more precise selection of parameters to freeze, one can employ the following alternative—and computationally more demanding—estimation method. We perform a binary search to identify the largest number $M$ of parameters that can be frozen without exceeding the error threshold. At each iteration $k$, we consider deactivating the $M_{k}$ least relevant parameters in our ordered list. We then solve the tVMC equations (3), restricting the quantum geometric tensor $S$ and force vector $F$ to the subset of parameters that remain active. This allows us to estimate the resulting $\varepsilon^{2}(t)$ when freezing those $M_{k}$ parameters. If this estimate exceeds the threshold $\lambda^{2}_{\text{LITE}}$, we decrease $M_{k}$; otherwise, we increase it. The binary search proceeds until the maximal number of parameters that can be safely deactivated—while maintaining the error below the threshold—is found.
@@ II.3.4 Preventing overparameterization
In certain situations, the expressivity of the ansatz may be insufficient to keep the squared LITE below the threshold $\lambda^{2}_{\text{LITE}}$, even when all the variational parameters are active. However, not all of these parameters may contribute meaningfully to the dynamics; some may correspond to redundant or physically irrelevant degrees of freedom. The inclusion of such parameters in the evolution can induce numerical instabilities caused by overparameterization [37].
@@ III Results
To evaluate the effectiveness of our atVMC algorithm, we benchmark it on the one-dimensional transverse-field Ising model (TFI) with periodic boundary conditions. This model, which is exactly solvable [53], is defined by the Hamiltonian
@@
where the index $i$ labels the spin sites, and $\hat{\sigma}_{i}^{x}$ and $\hat{\sigma}_{i}^{z}$ are Pauli operators. The model is characterized by the dimensionless parameter $g=|h/J|$ and exhibits a quantum critical point at $g=g_{c}\equiv 1$, which marks a quantum phase transition: for $g>g_{c}$ the ground state is paramagnetic, while for $g<g_{c}$ it is ferromagnetic. To explore nontrivial dynamics, we perform tVMC simulations of quantum quenches. In a quantum quench, the system is initially prepared in the ground state corresponding to some value $g=g_{1}$ (the ground state in our simulations is variationally optimized via stochastic reconfiguration [54, 55]), and then evolved with a different parameter $g=g_{2}$. We refer to this protocol as the quench $g=(g_{1}\to g_{2})$. Simulating quenches where $g_{2}$ is near the critical point $g_{c}$ is particularly challenging, due to the expected divergence of correlations [43].
We test atVMC on two commonly employed variational wave functions. The first one is the spin-Jastrow ansatz [56] with complex-valued parameters, defined as
@@
The second variational ansatz we consider is the restricted Boltzmann machine (RBM) wave function as introduced in Ref. [17], which is defined as
@@
All simulation results presented in this work are freely available online [57].
@@ III.1 Spin-Jastrow wave function, quench $g=(4\to 2)$
To illustrate the behavior of the atVMC algorithm, we begin by considering the quench $g=(4\to 2)$ in a TFI chain of $32$ spins. The system starts in a paramagnetic initial state and evolves under a different but still paramagnetic Hamiltonian. We employ the spin-Jastrow variational wave function, which in this case has $16$ independent parameters, and impose a LITE threshold of $\lambda^{2}_{\text{LITE}}=10^{-2}\cdot\operatorname{Var}(\hat{H})/\hbar^{2}$. The tVMC equations of motion are solved using the pseudoinverse regularization with a tolerance of $10^{-7}$ [17, 43, 38, 37]. Here, we employ the atVMC algorithm in its simplest form, allowing at most one parameter to be frozen or unfrozen per time step.
Fig. 1(a) shows the time evolution of the magnetization along the $x$-axis, defined as $\sigma_{x}=\left\langle\sum_{i}\hat{\sigma}^{x}_{i}\right\rangle/N$, and compares the variational and exact solutions, which are in good agreement.
Fig. 1(b) shows the number of active parameters as a function of time. All parameters are initially active, but during the first steps nearly all are quickly frozen: quantum correlations generated by the quench spread locally at first, and only the parameter encoding nearest-neighbor correlations needs to remain active. As the correlations propagate further through the system, the parameter describing next-nearest-neighbor correlations is the first to be unfrozen, followed by those corresponding to distance three, then four, and so on. At longer times, the pattern becomes more complex, with some medium-range parameters intermittently frozen.
Fig. 1(c) plots $\varepsilon^{2}(t)$, i.e., the squared LITE over time. Comparing it with panel (b), we observe that whenever $\varepsilon^{2}$ exceeds the threshold $\lambda^{2}_{\text{LITE}}$, a parameter is unfrozen, resulting in a sudden drop in the error. Conversely, when $\varepsilon^{2}$ is below $\lambda^{2}_{\text{LITE}}$, the algorithm freezes a parameter if doing so keeps the LITE within bounds. In these cases, $\varepsilon^{2}$ exhibits a sharp rise.
Figure 1: Quench $g=(4\to 2)$ simulated using atVMC with a spin-Jastrow ansatz on a TFI chain of $32$ spins. (a) Transverse magnetization per spin: variational results (solid line) compared with the exact solution (dashed line). Error bars are omitted as they are comparable in magnitude to the observable’s oscillations. (b) Number of active variational parameters selected by the adaptive algorithm. (c) Squared LITE during the evolution (solid line), compared to the target threshold (horizontal dashed line).
@@
To illustrate the importance of the collective parameter updates introduced in Sec. II.3.3, we now consider the quench $g=(4\to 1.5)$ in the $32$-spin TFI chain. This quench drives the system closer to the quantum critical point and generates more intricate quantum correlations, so we use a more expressive variational ansatz: a restricted Boltzmann machine (RBM) with density $d=3$, corresponding to $100$ variational parameters in our setup. We set the LITE threshold to $\lambda^{2}_{\text{LITE}}=10^{-3}\cdot\operatorname{Var}(\hat{H})/\hbar^{2}$ and employ pseudoinverse regularization with a tolerance of $10^{-7}$. In addition, we adopt the adaptive time-step scheme for the Heun integrator introduced in Ref. [38], which is compatible with atVMC upon a minor modification: the number of active parameters is updated only once every two steps.
Fig. 2 compares simulations with and without the collective parameter update. Panel (a) plots the evolution of the transverse magnetization $\sigma_{x}$, showing that the dynamics of both the simulations is in quite good agreement with the exact result. Panels (b) and (c) show the number of active parameters and the squared LITE as functions of time, respectively. As in the previous quench, the algorithm initially wants only a small subset of parameters to be active. At later times, however, all parameters are eventually unfrozen, since the full expressivity of the RBM with $d=3$ is insufficient to keep the error within the imposed bound.
@@
Figure 2: Quench $g=(4\to 1.5)$ simulated using atVMC with a RBM ansatz with density $d=3$ on a TFI chain of $32$ spins. Simulations are performed with an adaptive time-step scheme, both with (circular markers) and without (triangular markers) collective parameter updates. (a) Transverse magnetization per spin: atVMC results compared to the exact solution (dashed line). Error bars are omitted as they are smaller than the marker size. (b) Number of active variational parameters selected by the adaptive algorithm. (c) Squared LITE during the evolution, compared to the target threshold (horizontal dashed line).
@@
We now consider the quench $g=(0.5\to 1)$ in the 32-spin TFI chain. The initial ground state, prepared in the ferromagnetic phase, evolves under a Hamiltonian at the quantum critical point, so that simulating the quantum correlations generated by the quench is especially challenging [43]. In this case, we employ a particularly expressive ansatz, a RBM with density $d=15$, corresponding to $496$ variational parameters in our setup. We set the LITE threshold to $\lambda^{2}_{\text{LITE}}=10^{-3}\cdot\operatorname{Var}(\hat{H})/\hbar^{2}$, and employ pseudoinverse regularization with a tolerance of $10^{-7}$.
@@
The curve with circular markers in Fig. 3(a) shows the evolution of the transverse magnetization $\sigma_{x}$ for an atVMC simulation that employs collective parameter updates but no mechanism to prevent overparameterization. In this case, the adaptive adjustment of the number of active parameters is insufficient to avoid instabilities. This is due to the difficulty of the quench: after a certain point, even the full expressivity of the ansatz cannot keep the LITE below the threshold. As shown in Fig. 3(b), the algorithm responds by activating all available parameters in an attempt to reduce the error. However, this also reintroduces many irrelevant parameters, which contributes to the onset of instabilities.
@@
This issue is addressed by the overparameterization prevention scheme described in Sec. II.3.4. The curve with triangular markers in Fig. 3(a) shows the evolution of $\sigma_{x}$ for an atVMC simulation that incorporates this control mechanism with significance threshold $\eta_{\text{sig}}^{2}=5\cdot 10^{-3}$, which successfully prevents instabilities. The transverse magnetization remains compatible with the exact dynamics up to a certain point in the simulation, beyond which deviations arise due to the challenging nature of this quench. As shown in Fig. 3(b), even when the LITE exceeds the threshold, not all parameters are activated—an outcome of the overparameterization control scheme. Fig. 3(c) compares the squared LITE for the two simulations—with and without this control mechanism—and shows that, while the control scheme yields a slightly higher LITE, it remains of the same order of magnitude, while achieving substantially improved stability.
@@
Figure 3: Quench $g=(0.5\to 1)$ simulated using atVMC with a RBM ansatz with density $d=15$ on a TFI chain of $32$ spins, with collective parameter updates, with (triangular markers) and without (circular markers) overparameterization control. (a) Transverse magnetization per spin: atVMC results compared to the exact solution (dashed line). Error bars are omitted as they are comparable in magnitude to the observable’s oscillations. (b) Number of active variational parameters selected by the adaptive algorithm. (c) Squared LITE during the evolution, compared to the target threshold (horizontal dashed line).
@@
We now compare the performance of atVMC with that of the standard tVMC algorithm using commonly adopted regularization strategies, employing the same number of Monte Carlo samples per time step (here we use about $7\cdot 10^{4}$ Monte Carlo samples per time step). Fig. 4 contrasts the atVMC results with those obtained from a standard tVMC simulation that employs an adaptive time-step scheme combined with the signal-to-noise ratio (SNR) regularization introduced in Ref. [38]. The SNR threshold is set to $4$, a value specifically tuned to prevent instabilities, with energy conservation serving as our primary diagnostic. For the atVMC simulations, we use pseudoinverse regularization with a tolerance of $10^{-7}$. We note in passing that applying this same pseudoinverse regularization in standard tVMC leads to instabilities for this particular case.
@@
As a final example, we illustrate how the atVMC framework can be employed primarily as a regularization tool in this quench scenario. This is achieved by setting $\lambda_{\text{LITE}}^{2}$ to a very low value, close to the minimum LITE attainable by the ansatz during the simulation. The square-marked curves in Fig. 4 correspond to $\eta_{\text{sig}}^{2}=10^{-2}$ and $\lambda^{2}_{\text{LITE}}=10^{-5}\cdot\operatorname{Var}(\hat{H})/\hbar^{2}$. As shown in Panel (b), the full expressivity of the RBM with density $d=15$ is barely sufficient to reach this error threshold, and the LITE remains above it for most of the simulation. Nevertheless, choosing a threshold that is not strictly unreachable ensures that a collective parameter update is triggered at the beginning of the simulation, allowing the algorithm to identify an initial subset of relevant parameters. Had the threshold been too low to reach even momentarily, no collective freezing would have satisfied the accuracy condition, and all parameters would have remained active throughout. During the rest of the evolution, since the LITE generally stays above the threshold, the number of active parameters is primarily controlled by the overparameterization prevention scheme, effectively enforcing the maximal expressivity compatible with excluding non-relevant parameters, where relevance is defined by the significance threshold $\eta_{\text{sig}}^{2}$.
@@
Figure 4: Quench $g=(0.5\to 1)$ simulated using a RBM ansatz with density $d=15$ on a TFI chain of 32 spins. Three approaches are compared: standard tVMC with SNR regularization (circular markers), atVMC with $\lambda^{2}_{\text{LITE}}=10^{-3}\cdot\operatorname{Var}(\hat{H})/\hbar^{2}$ and $\eta_{\text{sig}}^{2}=5\cdot 10^{-3}$ (triangular markers), and atVMC with $\lambda^{2}_{\text{LITE}}=10^{-5}\cdot\operatorname{Var}(\hat{H})/\hbar^{2}$ and $\eta_{\text{sig}}^{2}=10^{-2}$ (square markers). (a) Transverse magnetization per spin: variational results compared to the exact solution (dashed line). Error bars are omitted as they are comparable in magnitude to the observable’s oscillations. (b) Squared LITE during the evolution, compared to the target thresholds $\lambda^{2}_{\text{LITE}}=10^{-3}\cdot\operatorname{Var}(\hat{H})/\hbar^{2}$ (horizontal dashed line) and $\lambda^{2}_{\text{LITE}}=10^{-5}\cdot\operatorname{Var}(\hat{H})/\hbar^{2}$ (horizontal dash-dotted line).

### v5

@@ II.3.3 Collective parameter update
For situations that demand a more precise selection of parameters to freeze, one can employ the following alternative—and computationally more demanding—estimation method. A binary search is performed to identify the largest number $M$ of parameters that can be frozen without exceeding the error threshold. At each iteration $k$, the $M_{k}$ least relevant parameters in our ordered list are considered for reactivation. Then, the tVMC equations (3) are solved, restricting the quantum geometric tensor $S$ and force vector $F$ to the subset of parameters that remain active. This procedure allows estimation of the resulting $\varepsilon^{2}(t)$ when those $M_{k}$ parameters are frozen. If this estimate exceeds the threshold $\lambda^{2}_{\text{LITE}}$, $M_{k}$ is decreased; otherwise, it is increased. The binary search proceeds until the maximal number of parameters that can be safely deactivated—while maintaining the error below the threshold—is found.
@@ II.3.4 Preventing overparameterization
In certain situations, the expressivity of the ansatz may be insufficient to keep the squared LITE below the threshold $\lambda^{2}_{\text{LITE}}$, even when all the variational parameters are active. However, not all of these parameters may contribute meaningfully to the dynamics; some may correspond to redundant or physically irrelevant degrees of freedom. The inclusion of such parameters in the evolution can induce numerical instabilities caused by overparameterization [Hofmann_Role_of_stochastic_noise].
@@ III Results
To evaluate the effectiveness of our atVMC algorithm, we benchmark it on the one-dimensional transverse-field Ising model (TFI) with periodic boundary conditions. This model, which is exactly solvable [PFEUTY1970], is defined by the Hamiltonian
@@
where the index $i$ labels the spin sites, and $\hat{\sigma}_{i}^{x}$ and $\hat{\sigma}_{i}^{z}$ are Pauli operators. The model is characterized by the dimensionless parameter $g=|h/J|$ and exhibits a quantum critical point at $g=g_{c}\equiv 1$, which marks a quantum phase transition: for $g>g_{c}$ the ground state is paramagnetic, while for $g<g_{c}$ it is ferromagnetic. To explore nontrivial dynamics, we perform tVMC simulations of quantum quenches. In a quantum quench, the system is initially prepared in the ground state corresponding to some value $g=g_{1}$ (the ground state in our simulations is variationally optimized via stochastic reconfiguration [Becca_Sorella_Optimization_Chapter, Sorella_stoc_recnfig]), and then evolved with a different parameter $g=g_{2}$. We refer to this protocol as the quench $g=(g_{1}\to g_{2})$. Simulating quenches where $g_{2}$ is near the critical point $g_{c}$ is particularly challenging, due to the expected growth of quantum correlations [quench_critical_point].
We test atVMC on two commonly employed variational wave functions. The first one is the spin-Jastrow ansatz [Spin_Jastrow_state] with complex-valued parameters, defined as
@@
The second variational ansatz we consider is the restricted Boltzmann machine (RBM) wave function as introduced in Ref. [Carleo_Solving_many-body_with_neural_networks], which is defined as
@@
All simulation results presented in this work are freely available online [salioni_2025_15736912].
@@ III.1 Spin-Jastrow wave function, quench $g=(4\to 2)$
To illustrate the behavior of the atVMC algorithm, we begin by considering the quench $g=(4\to 2)$ in a TFI chain of $32$ spins. The system starts in a paramagnetic initial state and evolves under a different but still paramagnetic Hamiltonian. We employ the spin-Jastrow variational wave function, which in this case has $16$ independent parameters, and impose a LITE threshold of $\lambda^{2}_{\text{LITE}}=10^{-2}\cdot\operatorname{Var}(\hat{H})/\hbar^{2}$. The tVMC equations of motion are solved using the pseudoinverse regularization with a tolerance of $10^{-7}$ [Carleo_Solving_many-body_with_neural_networks, quench_critical_point, Schmitt_and_Heyl, Hofmann_Role_of_stochastic_noise]. Here, we employ the atVMC algorithm in its simplest form, allowing at most one parameter to be frozen or unfrozen per time step.
Fig. 1(a) shows the time evolution of the magnetization along the $x$-axis, defined as $\sigma_{x}=\left\langle\sum_{i}\hat{\sigma}^{x}_{i}\right\rangle/N$. We compare the result obtained via atVMC with the exact solution and with the result obtained via standard tVMC, where all the parameters of the variational ansatz are active. The variational results are in very good agreement with the exact transverse magnetization and confirm that atVMC does not compromise accuracy when the ansatz is already well-conditioned. In order to better appreciate the level of agreement, Fig. 1(b) shows the difference of the transverse magnetization per spin computed with the Spin-Jastrow variational wave function and the exact value. Fig. 1(c) shows the number of active parameters as a function of time. All parameters are initially active, but during the first steps nearly all are quickly frozen: quantum correlations generated by the quench spread locally at first, and only the parameter encoding nearest-neighbor correlations needs to remain active. As the correlations propagate further through the system, the parameter describing next-nearest-neighbor correlations is the first to be unfrozen, followed by those corresponding to distance three, then four, and so on. At longer times, the pattern becomes more complex, with some medium-range parameters intermittently frozen. Fig. 1(d) plots $\varepsilon^{2}(t)$, i.e., the squared LITE over time. Comparing it with panel (c), we observe that whenever $\varepsilon^{2}$ exceeds the threshold $\lambda^{2}_{\text{LITE}}$, a parameter is unfrozen, resulting in a sudden drop in the error which, interestingly, aligns with the error of the tVMC simulation. Conversely, when $\varepsilon^{2}$ is below $\lambda^{2}_{\text{LITE}}$, the algorithm freezes a parameter if doing so keeps the LITE within bounds. In these cases, $\varepsilon^{2}$ exhibits a sharp rise.
Figure 1: Quench $g=(4\to 2)$ simulated using atVMC (circular markers, orange color) and tVMC (square markers, blue color) with a spin-Jastrow ansatz on a TFI chain of $32$ spins. (a) Transverse magnetization per spin: variational results compared with the exact solution (dashed line). Error bars are omitted as they are comparable in magnitude to the observable’s oscillations. (b) Difference between the variational and the exact transverse magnetization per spin. (b) Number of active variational parameters. (c) Squared LITE during the evolution, compared to the target threshold (horizontal dashed line).
@@
To illustrate the importance of the collective parameter updates introduced in Sec. II.3.3, we now consider the quench $g=(4\to 1.5)$ in the $32$-spin TFI chain. This quench drives the system closer to the quantum critical point and generates more intricate quantum correlations, so we use a more expressive variational ansatz: a restricted Boltzmann machine (RBM) with density $d=3$, corresponding to $100$ variational parameters in our setup. We set the LITE threshold to $\lambda^{2}_{\text{LITE}}=10^{-3}\cdot\operatorname{Var}(\hat{H})/\hbar^{2}$ and employ pseudoinverse regularization with a tolerance of $10^{-7}$. In addition, we adopt the adaptive time-step scheme for the Heun integrator introduced in Ref. [Schmitt_and_Heyl], which is compatible with atVMC upon a minor modification: the number of active parameters is updated only once every two steps.
Fig. 2 compares atVMC simulations—with and without the collective parameter update—and a standard tVMC simulation, in which all the variational parameters of the ansatz are evolved. Panel (a) plots the evolution of the transverse magnetization $\sigma_{x}$, showing that the dynamics of the atVMC simulations is in quite good agreement with the exact result and the tVMC simulation. Panels (b) and (c) show the number of active parameters and the squared LITE as functions of time, respectively. As in the previous quench, the algorithm initially wants only a small subset of parameters to be active. At later times, however, all parameters are eventually unfrozen, since the full expressivity of the RBM with $d=3$ is insufficient to keep the error within the imposed bound.
@@
Figure 2: Quench $g=(4\to 1.5)$ simulated using atVMC and tVMC with a RBM ansatz with density $d=3$ on a TFI chain of $32$ spins. atVMC simulations are performed with an adaptive time-step scheme, both with (diamond markers, red color) and without (circular markers, orange color) collective parameter updates. (a) Transverse magnetization per spin: variational results compared to the exact solution (dashed line). Error bars are omitted as they are smaller than the marker size. (b) Number of active variational parameters. (c) Squared LITE during the evolution, compared to the target threshold (horizontal dashed line).
@@
We now consider the quench $g=(0.5\to 1)$ in the 32-spin TFI chain. The initial ground state, prepared in the ferromagnetic phase, evolves under a Hamiltonian at the quantum critical point, so that simulating the quantum correlations generated by the quench is especially challenging [quench_critical_point]. In this case, we employ a particularly expressive ansatz, a RBM with density $d=15$, corresponding to $496$ variational parameters in our setup. We set the LITE threshold to $\lambda^{2}_{\text{LITE}}=10^{-3}\cdot\operatorname{Var}(\hat{H})/\hbar^{2}$, and employ pseudoinverse regularization with a tolerance of $10^{-7}$.
@@
The curve with circular markers in Fig. 3(b) shows the evolution of the transverse magnetization $\sigma_{x}$ for an atVMC simulation that employs collective parameter updates but no mechanism to prevent overparameterization. In this case, the adaptive adjustment of the number of active parameters is insufficient to avoid instabilities. Such instabilities are clearly visible in Fig. 3(a) where the evolution of the energy per spin is shown: in any stable tVMC or atVMC simulation, the energy is conserved throughout the evolution; thus, the presence of discontinuities, as observed here, provides a clear signature of numerical instabilities. This is due to the difficulty of the quench: after a certain time, even the full expressivity of the ansatz cannot keep the LITE below the threshold. As shown in Fig. 3(c), the algorithm responds by activating all available parameters in an attempt to reduce the error. However, this also reintroduces many irrelevant parameters, which contributes to the onset of instabilities.
@@
This issue is addressed by the overparameterization prevention scheme described in Sec. II.3.4. The curve with triangular markers in Fig. 3(b) shows the evolution of $\sigma_{x}$ for an atVMC simulation that incorporates this control mechanism with significance threshold $\eta_{\text{sig}}^{2}=5\cdot 10^{-3}$, which successfully prevents instabilities, as further confirmed by the good energy conservation shown in Fig. 3(a). The transverse magnetization remains compatible with the exact dynamics up to a certain point in the simulation, beyond which deviations arise due to the challenging nature of this quench. As shown in Fig. 3(c), even when the LITE exceeds the threshold, not all parameters are activated—an outcome of the overparameterization control scheme. Fig. 3(d) compares the squared LITE for the two simulations—with and without this control mechanism—and shows that, while the control scheme yields a slightly higher LITE, it remains of the same order of magnitude, while achieving substantially improved stability.
@@
Figure 3: Quench $g=(0.5\to 1)$ simulated using atVMC with a RBM ansatz with density $d=15$ on a TFI chain of $32$ spins, with collective parameter updates, with (diamond markers, red color) and without (circular markers, orange color) overparameterization control. (a) Energy per spin. (b) Transverse magnetization per spin: atVMC results compared to the exact solution (dashed line). Error bars are omitted as they are comparable in magnitude to the observable’s oscillations. (c) Number of active variational parameters selected by the adaptive algorithm. (d) Squared LITE during the evolution, compared to the target threshold (horizontal dashed line).
@@
We now compare the performance of atVMC with that of the standard tVMC algorithm using commonly adopted regularization strategies, employing the same number of Monte Carlo samples per time step (here we use about $7\cdot 10^{4}$ Monte Carlo samples per time step). Fig. 4 contrasts the atVMC results with those obtained from a standard tVMC simulation that employs an adaptive time-step scheme combined with the signal-to-noise ratio (SNR) regularization introduced in Ref. [Schmitt_and_Heyl]. The SNR threshold is set to $4$, a value specifically tuned to prevent instabilities, with energy conservation serving as our primary diagnostic. For the atVMC simulations, we use pseudoinverse regularization with a tolerance of $10^{-7}$. We note in passing that applying this same pseudoinverse regularization in standard tVMC leads to instabilities for this particular case.
@@
In the next example, we illustrate how the atVMC framework can be employed primarily as a regularization tool in this quench scenario. This is achieved by setting $\lambda_{\text{LITE}}^{2}$ to a very low value, close to the minimum LITE attainable by the ansatz during the simulation. The square-marked curves in Fig. 4 correspond to $\eta_{\text{sig}}^{2}=10^{-2}$ and $\lambda^{2}_{\text{LITE}}=10^{-5}\cdot\operatorname{Var}(\hat{H})/\hbar^{2}$. As shown in Panel (b), the full expressivity of the RBM with density $d=15$ is barely sufficient to reach this error threshold, and the LITE remains above it for most of the simulation. Nevertheless, choosing a threshold that is not strictly unreachable ensures that a collective parameter update is triggered at the beginning of the simulation, allowing the algorithm to identify an initial subset of relevant parameters. Had the threshold been too low to reach even momentarily, no collective freezing would have satisfied the accuracy condition, and all parameters would have remained active throughout. During the rest of the evolution, since the LITE generally stays above the threshold, the number of active parameters is primarily controlled by the overparameterization prevention scheme, effectively enforcing the maximal expressivity compatible with excluding non-relevant parameters, where relevance is defined by the significance threshold $\eta_{\text{sig}}^{2}$.
@@
Figure 4: Quench $g=(0.5\to 1)$ simulated using a RBM ansatz with density $d=15$ on a TFI chain of 32 spins. Three approaches are compared: standard tVMC with SNR regularization and an adaptive time-step scheme (square markers, blue color), atVMC with $\lambda^{2}_{\text{LITE}}=10^{-3}\cdot\operatorname{Var}(\hat{H})/\hbar^{2}$ and $\eta_{\text{sig}}^{2}=5\cdot 10^{-3}$ (circular markers, orange color), and atVMC with $\lambda^{2}_{\text{LITE}}=10^{-5}\cdot\operatorname{Var}(\hat{H})/\hbar^{2}$ and $\eta_{\text{sig}}^{2}=10^{-2}$ (diamond markers, red color). (a) Transverse magnetization per spin: variational results compared to the exact solution (dashed line). Error bars are omitted as they are comparable in magnitude to the observable’s oscillations. (b) Squared LITE during the evolution, compared to the target thresholds $\lambda^{2}_{\text{LITE}}=10^{-3}\cdot\operatorname{Var}(\hat{H})/\hbar^{2}$ (horizontal dashed line) and $\lambda^{2}_{\text{LITE}}=10^{-5}\cdot\operatorname{Var}(\hat{H})/\hbar^{2}$ (horizontal dash-dotted line).
III.4 RBM wave function, quench $g=(4\to 2)$
Figure 5: Quench $g=(4\to 2)$ simulated using atVMC with a RBM ansatz with density $d=1$ on a TFI chain of $128$ spins. Two approaches are compared: standard tVMC (square markers, blue color) and atVMC with $\lambda^{2}_{\text{LITE}}=10^{-4}\cdot\operatorname{Var}(\hat{H})/\hbar^{2}$ and $\eta_{\text{sig}}^{2}=10^{-2}$ (circular markers, orange color). (a) Energy per spin. (b) Transverse magnetization per spin: atVMC and tVMC results compared to the exact solution (dashed line). Error bars are omitted as they are comparable in magnitude to the observable’s oscillations. (c) Difference between the variational and the exact transverse magnetization per spin for the atVMC simulation and the tVMC simuation. (d) Number of active variational parameters selected by the adaptive algorithm. (e) Squared LITE during the evolution, compared to the target threshold (horizontal dashed line).
As a final example, we reconsider the quench $g=(4\to 2)$ in a TFI chain of $128$ spins. Increasing the system size also increases the number of variational parameters, so even this relatively simple quench can enter a regime of potential overparameterization. We use a restricted Boltzmann machine (RBM) with density $d=1$, corresponding to 130 variational parameters in our setup. We set the LITE threshold to $\lambda^{2}_{\text{LITE}}=10^{-4}\cdot\operatorname{Var}(\hat{H})/\hbar^{2}$ and the significance threshold to $\eta_{\text{sig}}^{2}=10^{-2}$. The pseudoinverse regularization is applied with a tolerance of $10^{-7}$. We compare the atVMC results with a tVMC simulation using the same regularization scheme with identical tolerance.
Fig. 5(a) shows the time evolution of the energy per spin. In the standard tVMC simulation, the applied regularization scheme is not sufficient to suppress minor numerical instabilities, as evidenced by the small jump in Fig. 5(a) around $t\simeq 0.25\,\hbar/J$. Nevertheless, the atVMC simulation remains stable under the same conditions. Figures 5(b) and (c) demonstrate that the variational results are in excellent agreement with the exact evolution, with panel (b) showing the transverse magnetization per spin and panel (c) displaying the corresponding deviations. A small jump in the transverse magnetization obtained from the standard tVMC simulation can be also detected at $t\simeq 0.25\,\hbar/J$, coinciding with the numerical instability observed for the energy. In Fig. 5(d) the substantial reduction in the number of evolved variational parameters during the atVMC simulation is evident, while panel (e) shows that the squared LITE remains largely unaffected.